The summit of Mount Everest is 8848.0 m above sea
 level, making it the highest summit on Earth. In 1953, Edmund Hillary was the first person to reach the summit. Suppose upon reaching there, Hillary slid a rock from rest with a 32.0 g mass down the side of the mountain. If the rock's speed was 24.0 m/s when it was 7690.0 m above sea level, how much work was done
on the rock by air resistance?







1
Expert's answer
2022-01-27T13:17:20-0500

"m=32.0 \;g"

"h_1=8848.0 \;m"

"h_2=7690.0\;m"

"v_1=0 \;m/s"

"v_2=24.0 \;m/s"

"g=9.81 \;m/s^2"


"W_r=m(\frac{1}2(v_2^2-v_1^2)-g(h_1-h_2))"

"W_r=(32\times 10^{-3} kg )(\frac{1}2(24^2-0^2) m^2/s^2 -9.81m/s^2(8848-7690)m)"

"W_r=(32\times 10^{-3}kg)(288-11358.4)m^2/s^2"

"W_r(32\times 10^{-3}kg)(-11060.4m^2/s^2)=-353.9J"


Answer: "W_r=-353.9 \;J"